Saturday 28 August 1999 3.00am EST
|The following is an overview of the events that occurred over the last 24 hours.
The P8 load continued to execute as expected. A segmented maneuver scheduled for 10:28-11:33pm on 8/27 completed successfully and confirmed confirmed that the corrective actions taken after the safe mode entry were effective. The maneuver was performed during a communications pass and with appropriate engineering staff on console. The maneuver took Chandra to the bright star Capella in preparation for the deployment of the High Energy Transmission Grating (HETG).
The HETG was opened successfully at 1:35 am on 8/28. The procedure opened both launch locks and commanded the grating to the inserted position via hardware commands. All engineering data appeared nominal during the grating insertion. Data received by ACIS-S within 15 minutes showed the anticipated "X" pattern from the HEG and MEG facets and indicated that the HETG was inserted and operating as expected. A short time later features were evident in the spectrum. The cheer from the science room was audible in the control room! The grating will remain in place for 40 hours to obtain calibration data with ACIS-S then be retracted.
Return to Status Reports